The speed of a boat in still water is 10 km/h. The boat travels 12 km upstream and 28 km downstream in a total time of 4 hours. What is the speed of the stream?
Answer by jorel1380(2518)

12/(10-x)+28/(10+x)=4
12(10+x)+28(10-x)=400-4x2
120+12x+280-28x=400-4x2
4x2-16x=0
x2-4x=0
x2=4x
x=4
The speed of the stream is 4 kph..